# Topological Indices With Degree Sequence $\mathscr{D}$ of Tree

### v1

@@ 3.2. Sigma Index With Degree Sequence $d_{n}\geqslant\dots\geqslant d_{1}$
Example 1.
Let be a degree sequence is $d=(4,8,10,14,18,20)$, then for Max sigma we have:
$[14802,18,4,10,14,8,20],[14802,18,4,10,14,8,20],[14802,18,4,14,8,10,20],[14802,18,4,14,10,8,20],$
$[14802,18,8,10,14,4,20],[14802,18,8,14,4,10,20],[14802,18,8,14,10,4,20],[14802,18,10,4,14,8,20],$
$[14802,18,10,8,14,4,20],[14802,20,4,10,14,8,18],[14802,20,4,14,8,10,18],[14802,20,4,14,10,8,18],$
$[14802,20,8,10,14,4,18],[14802,20,8,14,4,10,18],[14802,20,8,14,10,4,18],[14802,20,10,4,14,8,18],$
$[14802,20,10,8,14,4,18].$
and for Min sigma we have:
$[14196,4,10,14,18,20,8],[14196,4,10,14,18,20,8],[14196,4,10,14,20,18,8],[14196,4,10,18,20,14,8],$
$[14196,4,10,20,18,14,8],[14196,4,14,18,20,10,8],[14196,4,14,20,18,10,8],[14196,4,18,20,14,10,8],$
$[14196,4,20,18,14,10,8],[14196,8,10,14,18,20,4],[14196,8,10,14,20,18,4],[14196,8,10,18,20,14,4],$
$[14196,8,10,20,18,14,4],[14196,8,14,18,20,10,4],[14196,8,14,20,18,10,4],[14196,8,18,20,14,10,4],$
$[14196,8,20,18,14,10,4].$
@@ 4. Conclusion
Through this paper, we presented a study of the topological indices defined by the Sigma Index and the Albertson Index according to a degree sequence $d_{n}\geqslant\dots\geqslant d_{1}$, in Hypothesize 3 for a degree sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$ where $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then we defined Sigma index as $\sigma(T)=\sum_{i=1}^{n-2}d_{i}(d_{i+1})^{2}+(d_{1}-1)^{3}+(d_{6}-1)^{3}$. The examination of topological indices on trees gives us a general overview through bounds to find the maximum and minimum bounds and this is what we were keen to show in the proofs we presented for each case discussed.
Acknowledgments
I would like to extend my deepest gratitude to my thesis advisor, Dr. Artem Kurnosov, for her invaluable guidance and support throughout this research project. I also appreciate the insightful feedback provided by Prof.Belov Alexey Yakovlevich during the review process.

### v2

@@ 4. Conclusion
Through this paper, we presented a study of the topological indices defined by the Sigma Index and the Albertson Index according to a degree sequence $d_{n}\geqslant\dots\geqslant d_{1}$, in Hypothesize 3 for a degree sequence $\mathscr{D}=(d_{1},\dots,d_{6})$ where $d_{6}\geqslant d_{5}\geqslant\dots\geqslant d_{1}$, then we defined Sigma index as $\sigma(T)=\sum_{i=1}^{n-2}d_{i}(d_{i+1})^{2}+(d_{1}-1)^{3}+(d_{6}-1)^{3}$. The examination of topological indices on trees gives us a general overview through bounds to find the maximum and minimum bounds and this is what we were keen to show in the proofs we presented for each case discussed.